OpenStudy (anonymous):

Find the distance between these points. R(-1, 0), S(8, 6) sqrt (26) sqrt (85) 3 sqrt (13)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

sqrt((-1-8)^2+(0-6)^2)=sqrt(81+36)=sqrt117=3*sqrt13 medal?

OpenStudy (alexandervonhumboldt2):

3 sqrt 13

OpenStudy (anonymous):

thnx